OpenStudy (anonymous):

How do I solve this expression:

OpenStudy (anonymous):

(x-12)(x^2 – 2x + 16)

OpenStudy (strawberry17):

You multiply x and -12 by each term in the second parentheses. Then combine like terms :)

OpenStudy (strawberry17):

Do you know how to do that?

OpenStudy (anonymous):

Well I can try hold on lemme get paper xD

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

Okay idk if i did this right but so far I got... 12x^2 - 24x + 192

OpenStudy (strawberry17):

You have some figures correct... when you multiply x by x^2 you will get x^3. Then you multiply x by -2x and 16. Then you will multiply -12 by x^2, -2x, and 16. Then combine like terms.

OpenStudy (anonymous):

okay

OpenStudy (strawberry17):

Let me know when you get it :)

OpenStudy (anonymous):

I think I just confused myself O.o Can u draw it out?

OpenStudy (strawberry17):

x • x^2 = x^3 x • -2x = -2x^2 x • 16 = 16x -12 • x^2 = -12x^2 -12 • -2x = 24x -12 • 16 = -192 Then combine like terms :)

OpenStudy (strawberry17):

x^3 - 2x^2 - 12x^2 + 16x + 24x - 192

OpenStudy (anonymous):

thanks for writing it out like that (im more of a visual learner) lol

OpenStudy (strawberry17):

ur welcome :)

OpenStudy (strawberry17):

i am too